OpenStudy (anonymous):

how do you find the domain of the function: u (x)= square root of -6x+12

OpenStudy (akashdeepdeb):

\(u(x) = \sqrt{-6x + 12}\) We know that the only restriction on the function is that: \(-6x+12 \ge 0\) Because for all real \(\sqrt{a} \implies a \ge 0\) Now, solve the inequality and you'll get the restrictions on the domain.